OpenStudy (anonymous):

PLS HELP WILL MEDAL simply square root of 36

OpenStudy (anonymous):

i think that is what this is asking

OpenStudy (trojanpoem):

6

OpenStudy (anonymous):

The square root of 36 is 6

OpenStudy (trojanpoem):

6 and - 6

OpenStudy (legends):

@candiimonstaa you are right because 6*6 equals to 36

OpenStudy (anonymous):

can you show me how you figure it out? bc this one is different

OpenStudy (trojanpoem):

I will try 36 = 6 * 6 or -6 * -6

OpenStudy (trojanpoem):

But if you have number like this 8 when you try to get it's root you won't find clear one so It will be root ( 4 * 2) the root of 4 is 2 so it will be 2rootof(2)

OpenStudy (legends):

@TrojanPoem ur right nice job

OpenStudy (trojanpoem):

Thanks.

OpenStudy (anonymous):

can u help me with the second attachment i sent

OpenStudy (trojanpoem):

Sure.

OpenStudy (anonymous):

thank u

OpenStudy (trojanpoem):

Instead of type root i will type r() r(2) * r(800) = r(800 * 2) = r(1600) which is r(100) * r(16) r(100) = 10 and r(16) =4 so it's 4 * 10 = 40

OpenStudy (trojanpoem):

typing*

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (anonymous):

ill try to do the last two, they're just hard for me

OpenStudy (trojanpoem):

If you failed, tell me i will help.

OpenStudy (anonymous):

it looks like this- they keep making them different

OpenStudy (anonymous):

is it 3/2

OpenStudy (trojanpoem):

Yeah

OpenStudy (anonymous):

not -3/2?

OpenStudy (trojanpoem):

Both are right , at least to me

OpenStudy (anonymous):

how do i figure out if its negative or not

OpenStudy (trojanpoem):

Always negative and positive until he says something else. for ex 36 is -6 * -6 = 36 and 6 * 6

OpenStudy (anonymous):

but they are both listed as an answer, but i can only choose one

OpenStudy (trojanpoem):

In this case choose the positive one.

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

is it c

OpenStudy (trojanpoem):

Sorry , was afk.

OpenStudy (trojanpoem):

Just sum them

OpenStudy (trojanpoem):

4r(7)

OpenStudy (anonymous):

thank you so much for ALL the help, genuinely appreciate it

OpenStudy (anonymous):

if i could give 3286238 medals to you, i would :)

OpenStudy (trojanpoem):

Thanks :D

OpenStudy (trojanpoem):

Enough for me to see you happy.